(4z + x²z) dz + (1+z²)dx - (yz² + y)(4+x²)dy=0


1
Expert's answer
2021-09-05T15:19:25-0400

By Lagrange method,

"\frac{dx}{1+z^2}=\frac{dy}{- (yz² + y)(4+x²)}=\frac{dz}{4z+x^2z}\\"

By taking the first equation,

"\frac{dy}{- (yz² + y)(4+x²)}=\frac{dz}{4z+x^2z}\\
\frac{dy}{- y(z² + 1)(4+x²)}=\frac{dz}{z(4+x^2)}\\
\frac{dy}{- y}=(\frac{1}{z}+z)dz\\"

Integrating both sides,

"-lny=lnz+\frac{z^2}{2}+c\\
ln(\frac{1}{yz})-\frac{z^2}{2}=c"

And, from the second equation,

"\frac{dx}{1+z^2}=\frac{dz}{4z+x^2z}\\
\frac{dx}{1+z^2}=\frac{dz}{z(4+x^2)}\\
(4+x^2)dx=(\frac{1}{z}+z)dz\\
4x+\frac{x^3}{3}=lnz+\frac{z^2}{2}+c\\
4x+\frac{x^3}{3}-lnz-\frac{z^2}{2}=c"

Thus, the general solution is given by

"F(ln(\frac{1}{yz})-\frac{z^2}{2}, 4x+\frac{x^3}{3}-lnz-\frac{z^2}{2})=0"